OpenStudy (anonymous):

y+11=-4x 13=y-2x setting them equal

OpenStudy (anonymous):

find the values of y from both the equations and set them equal..hence you get the value of x

OpenStudy (anonymous):

could you work it out plz

OpenStudy (anonymous):

Since y=-4x-11 and y=2x+13 set -4x-11=2x+13 and solve

OpenStudy (anonymous):

both are solved for y so you have two expressions that should be equal bc they both equal y

OpenStudy (anonymous):

-4x -11= 2x+13 +4x +4x -11 = 6x + 13 -13 -13 -24=6x __ __ 6 6 x = -4

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yeah hope i do good on my test tomorow

OpenStudy (anonymous):

me too ; ) anything else you need help with?

OpenStudy (anonymous):

yeah one more question 7x-y=8 9x+y=24 using combination